OpenStudy (anonymous):

4/5 more than the product of 2/3 and x is 7/15 Solve for x

OpenStudy (anonymous):

4/5+2/3=12/15+10/15=22/15 22/15-7/15=13/15 13/15-7/15=6/15 x=6 or x=2,2/3